Barb bought a house with 20% down and the rest financed by a 30-year mortgage with monthly payments calculated at a nominal annual rate of interest 8.4% compounded monthly.

She notices that one-third of the way through the mortgage she will still owe 200,000.

Determine the purchase price of the house.

  1. 252,706

  2. 262,706

  3. 272,806

  4. 282,706

  5. 292,706
Answer #1

One third of the mortgage is 10 years,

So,

At the end of Year 10, Barb owes $200,000

Calculating Monthly Payment,

Using TVM Calculation,

PMT = [PV = 200,000, FV = 0, N = 240, I = 0.084/12]

PMT = $1,723.01

Calculating Loan Amount,

Using TVM Calculation,

PV = [FV = 0, PMT = -1,723.01, N = 360, I = 0.084/12]

PV = $226,164.98

Value of House = 226,164.98/(0.80)

Value of House = $282,706